OpenStudy (anonymous):

hi need help in boolean algebra. (A'+BC)(D'E+F)'+(D'E+F) i know (D'E+F)'+(D'E+F)=1, but (A'+BC)(D'E+F)' should be done 1st,......

OpenStudy (ash2326):

we have \[(A'+BC)(D'E+F)'+(D'E+F)\] we know that in Boolean Algebra \[X+YZ=(X+Y)(X+Z)\] we get \[(D'E+F+(D'E+F)')(D'E+F+A'+BC)\] so we get \[(1)(D'E+F+A'+BC)\] we get \[D'E+F+A'+BC\]

OpenStudy (anonymous):

ooh tyvm, do you mind if i ask another 1? W'XY+WXZ+WY'Z+W'Z = W'XY+XZ+Y'Z+W'Z and im stucked. lol or i shouldnt approach it like this?

OpenStudy (anonymous):

should i take out the z from the last 3 terms?